OpenStudy (anonymous):

Separate 114 into two parts such that when the larger is divided by the smaller, the quotient is 3 and the remainder is 18.

OpenStudy (anonymous):

I know. I just changed it :)

OpenStudy (anonymous):

90 and 24

OpenStudy (ybarrap):

114 = a + b a/b = 3 + 18/b a = 3b + 18 114 = 3b + 18 + b = 4b + 18 4b = 114 - 18 = 96 b = 96/4 = 24 a = 114 - 24 = 90 Check a/b = 90/24 = 3 3/4 = 3 18/24 a + b = 90 + 24 = 114 Ans, a = 24, b = 90

OpenStudy (anonymous):

thanks :D